Calculate the heat radiated into a room at by a filament lamp, with a surface area
of 3 × 10-6 m2, an emissivity of 0.92 and a temperature of 2,500K. In doing this
calculation you can neglect the heat radiated back by the room at the lamp – why is this?
1
Expert's answer
2020-12-06T17:32:16-0500
"Q=e\sigma T^4A\\
Q=(0.92)(5.67\cdot10^{-8})(2500)^4(3\cdot10^{-6})=6.1\ J"